A tourist takes a trip in a hot air balloon . The hot air balloon is moving vertically upwards at a constant speed of 44m/s. At the height of 20m she accidentally drops her camera over the side of the balloon's basket. Calculate the displacement covered by the camera after 2 seconds.


1
Expert's answer
2021-01-29T16:55:46-0500

Let's first find the position of the camera at t=2 seconds after its release from the kinematic equation:


"y(2.0\ s)=y_0 +v_0t-\dfrac{1}{2}gt^2,""y(2.0\ s)=20\ m+44\ \dfrac{m}{s}\cdot2\ s-\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(2\ s)^2=88.4\ m."

Finally, we can calculate the diisplacement covered by the camera after 2 seconds:


"y=y_f-y_i=88.4\ m-20\ m=68.4\ m."